# FedSOL: Stabilized Orthogonal Learning in Federated Learning

### v4

@@ 1 Introduction
Recent research suggests that the deviation of local learning in FL resembles Catastrophic Forgetting in Continual Learning (CL) [33, 57, 67]. In CL, fitting the model to a new task can alter parameters critical for previous tasks, thus negatively impacting their performance [53, 45, 65, 48]. Similarly in FL, local learning is prone to be overfitted on local datasets, causing the model to forget global knowledge not represented in the local distributions [33, 69]. Consequently, FL must navigate the balance between preserving previous global knowledge and acquiring new local knowledge during local learning.
Inspired by CL, we aim to resolve the conflict between two objectives through Orthogonal Learning[19, 41]. In CL, an effective strategy for preserving knowledge from previous tasks is to minimizing the new task’s loss using gradients that are orthogonal to the loss space of the old tasks [56, 9, 10, 76]. These orthogonal gradients enable the model to reduce the loss on the new task without negatively affecting performance on previous tasks. Our primary motivation is to adapt this strategy to FL by identifying gradients that are orthogonal to directions affecting the global knowledge while still enabling effective training on local datasets.
@@ 2.3 Proximal Gradient Projection
None (FedAvg) 59.80±${}_{\pm 1.12}$
L2 Distance 59.80±${}_{\pm 1.12}$ 56.35±${}_{\pm 2.85}$ (- 3.45)
KL-Divergence 60.31±${}_{\pm 2.07}$ 50.88±${}_{\pm 3.55}$ (- 9.43)
@@ 3.1 Preliminary: Overview of SAM
In the above equation, the inner maximization identifies a parameter perturbation $\epsilon$ that maximizes loss change within the $\rho$-ball neighborhood. This is practically approximated by a single re-scaled gradient step ${\boldsymbol{\epsilon}}^{*}$$=$$\rho{\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})/\|{\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})\|_{2}$. The outer minimization is then conducted by a base optimizer, such as SGD [50], taking the gradient ${\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w}+{\boldsymbol{\epsilon}}^{*})$ at the perturbed weights. SAM demonstrates an exceptional ability to perform well across different model structures [11, 77] and tasks [66, 1] with high generalization performance. In FL, applying SAM improves the generalization of each client’s local model [55, 6]. However, since this approach only addresses local objectives, its effectiveness in generalization is mostly confined to local data distributions [60, 61] and and still encounter conflicts with the proximal objective.
@@ 3.2 Adversarial Proximal Perturbation
In the above procedures, the model is updated using the gradient of the local loss $\mathcal{L}^{k}_{\mathrm{local}}$, while the proximal loss $\mathcal{L}_{{p}}^{k}$ plays only an implicit role in perturbing weights. We emphasize that unlike SAM, which employs the same loss for both weight perturbations and parameter updates, FedSOL distinguishes between the roles of these two types of losses to address the knowledge trade-off in FL. A more detailed comparison of these two distinct approaches in the FL context is discussed in Appendix E.
@@ 3.5 Theoretical Analysis
Meanwhile, Proposition 2 compares the change of local loss $\mathcal{L}^{k}_{\mathrm{local}}$ under FedSOL to its counterpart in FedAvg. This proposition suggests that, although FedSOL calculates the local gradient at perturbed weights, its impact on the local loss is almost identical to that of FedAvg. This implies that FedSOL effectively reduces local loss without significantly slowing down the learning process. As a result, FedSOL successfully mitigates the conflict between the proximal objective and the local objective. The detailed proofs of the propositions are provided in Appendix G.
@@ 4.2 Proximal Orthogonality of FedSOL
[…]
@@ 4.3 Performance on Data Heterogeneity
Table 2 presents a comparison between our approach, FedSOL, and other baselines. The results show that many recently proposed FL methods tend to underperform even when compared to the standard FedAvg baseline. A similar observation is also reported in [75, 33], which points out that many FL methods are sensitive to the learning scenarios. In contrast, FedSOL achieves state-of-the-art results in most cases, consistently outperforming FedAvg across all evaluated scenarios. Particularly, FedSOL shows remarkable improvement at high heterogeneity levels, such as in Sharding ($s$=2) and LDA ($\alpha$=0.05) scenarios. We further provide the learning curves in Appendix C, comparison to other weights perturbation strategies in Appendix F, and personalized FL performance in Appendix D.
@@
We examine the learning factors that influence FedSOL’s performance in Figure 4. FedSOL consistently surpasses FedAvg across various factors, achieving its best performance within the $\rho$ range between 0.5 and 2.0. Most notably, FedSOL’s gains increase as a smaller portion of clients participate in each round. For example, FedAvg’s performance significantly declines at a sampling ratio of 0.02, falling to near-random accuracy. However, FedSOL remains robust under such conditions.
@@
Vgg11 FedAvg 41.30±${}_{\pm 1.07}$ 50.02±${}_{\pm 4.25}$ 61.79±${}_{\pm 9.88}$
FedProx 40.45±${}_{\pm 1.41}$ 31.07±${}_{\pm 6.72}$ 63.47±${}_{\pm 2.68}$
FedNTD 60.55±${}_{\pm 2.14}$ 56.62±${}_{\pm 2.64}$ 69.82±${}_{\pm 2.27}$
FedSOL 56.39±${}_{\pm 1.40}$ 74.74±${}_{\pm 0.04}$ 78.38±${}_{\pm 1.12}$
Res18 FedAvg 49.92±${}_{\pm 0.62}$ 76.98±${}_{\pm 2.90}$ 57.91±${}_{\pm 1.27}$
FedProx 59.00±${}_{\pm 2.58}$ 82.09±${}_{\pm 2.35}$ 75.84±${}_{\pm 1.58}$
FedNTD 57.79±${}_{\pm 3.42}$ 78.50±${}_{\pm 0.18}$ 76.87±${}_{\pm 0.57}$
FedSOL 66.32±${}_{\pm 0.48}$ 85.97±${}_{\pm 0.04}$ 80.59±${}_{\pm 0.11}$
SL-ViT FedAvg 35.48±${}_{\pm 2.09}$ 53.94±${}_{\pm 5.17}$ 72.44±${}_{\pm 1.91}$
FedProx 38.73±${}_{\pm 1.23}$ 58.25±${}_{\pm 4.23}$ 74.10±${}_{\pm 1.23}$
FedNTD 47.59±${}_{\pm 2.84}$ 61.46±${}_{\pm 1.76}$ 71.65±${}_{\pm 1.71}$
FedSOL 47.95±${}_{\pm 1.51}$ 67.19±${}_{\pm 0.33}$ 77.96±${}_{\pm 0.47}$
@@ 4.4 Ablation Study
The results in Table 5 reveal that perturbing only the last classifier layer (Head in Table 5) is sufficient for FedSOL. The performance reaches as high as the full-model perturbation, yet the required computation is considerably lower. Interestingly, perturbing all layers except the classifier head (Body in Table 5) requires almost the same amount of computation but results in a drop in performance, underscoring the significance of the later layers.
@@ 6 Related Work
In CL, many approaches have adopted orthogonal learning, which align new task gradients orthogonal to old task loss spaces [13, 56, 19, 41]. This typically involves retaining data or gradients as memory [9, 10]. In FL, few recent attempts have applied similar strategies,[4, 42]. For example, FOT [4] uses a random Gaussian matrix with SVD, and GradMA [42] employs gradient memory and quadratic programming, both requiring substantial computational resources. In our work, we use the proximal objective to preserve knowledge and update using a local gradient orthogonal to it. However, we observe that directly opposing the proximal gradient can adversely affect performance, a finding that aligns with recent CL research about the drawbacks of direct gradient projection [76, 70]. To address this, we propose FedSOL, which implicitly finds the orthogonal local gradient.
@@ Appendix A Table of Notations
$k$
Index for clients ($k\in[K]$)
$g$
Index for global server
@@
$\mathcal{D}$
Whole dataset
$\mathcal{D}^{k}$
Local dataset of the $k$-th client
$\alpha$
Concentration parameter for the Dirichlet distribution
$s$
The number of shards per user
@@
$\beta,\mu$
Multiplicative coefficient for the proximal loss
$\gamma$
Learning rate
$\tau$
Temperature to be divided in the softmax probability distribution
$\rho$
Perturbation Strength for SAM-related algorithms
$\boldsymbol{\Lambda}$
Vector consists with scaling parameters for perturbation vector in SAM-related algorithms
@@
$\boldsymbol{w}_{g}$
Weight of the global server model
$\boldsymbol{w}_{k}$
Weight of the $k$-th client model
$\|\boldsymbol{w}_{g}-\boldsymbol{w}_{k}\|$
Collection of $L^{2}$-norm between server and client models, among all rounds.
@@
$\mathcal{L}^{k}_{\mathrm{local}}$
Local objective for the $k$-th client
$\mathcal{L}^{k}_{p}$
Proximal objective for the $k$-th client
@@ B.2 Datasets
  • CINIC-10 [12] (90,000 / 90,000): contains a combination of CIFAR and downsized ImageNet [14], which is compiled to serve as a bridge between the two datasets. The data is augmented using Random Cropping, Random Horizontal Flipping, and Normalization.
  • PathMNIST [71] (110,000 / 7,180): contains non-overlapping patches from Hematoxylin & Eosin stained colorectal cancer histology slide images. The data is augmented using Random Horizontal Flipping, and Normalization.
  • TissueMNIST [71] (189,106 / 47,280): contains microscope images of human kidney cortex cells, which are segmented from 3 reference tissue specimens. The data is augmented using Random Horizontal Flipping, and Normalization. The data is converted to 3-channel RGB images.
@@ Appendix D Personalized Performance
Appendix E Comparison to the Sharpness-Aware Optimization
E.1 SAM Optimization in FL
@@ E.2 Limitations of using SAM in FL
Appendix F Other Perturbation Strategies
@@
We exclude the strategies that obtaining proximal loss at the perturbed weights using the local loss gradient i.e., $*\mathcal{L}_{p}(\boldsymbol{w}_{k}+\boldsymbol{\epsilon}_{c}^{*}$), where $\boldsymbol{\epsilon}_{c}^{*}$ $=$ $\rho\,\frac{\boldsymbol{g}_{p}+\boldsymbol{g}_{l}}{\lVert\boldsymbol{g}_{p}+\boldsymbol{g}_{l}\rVert}$, as it leads the learning to diverge. The detailed formulation for each method is provided in Table 10 with its corresponding performance. The results in Table 10 demonstrates that utilizing the local loss gradient at weights perturbed by the proximal loss gradient ($A_{1}$ in Table 10) yields outperforms the other approaches. We suggest that our FedSOL is an effective way to integrate proximal restriction effect into SAM optimization in FL.
@@
Appendix G Proof of Proposition

### v6

# FedSOL: Stabilized Orthogonal Learning with Proximal Restrictions in Federated Learning

@@ 1 Introduction
Recent research suggests that such deviation in FL resembles Catastrophic Forgetting in Continual Learning (CL) [33, 57, 67]. In CL, fitting the model to a new task alters parameters critical for previous tasks, thus impairing their performance [53, 45, 65, 48]. Similarly in FL, local learning is prone to be overfitted on local datasets, causing the model to forget global knowledge not represented in the local distributions [33, 69]. Consequently, FL must navigate the balance between preserving previous global knowledge and acquiring new local knowledge during local learning. Inspired by CL, we aim to resolve the conflict between two objectives through Orthogonal Learning[19, 41]. In CL, an effective strategy for preserving previous knowledge is to minimizing the new task’s loss using gradients that are orthogonal to the loss space of the old tasks [56, 9, 10, 76]. These orthogonal gradients enable the model to reduce the loss on the new task without negatively affecting performance on previous tasks. Our primary motivation is to adapt this strategy to FL by identifying gradients that are orthogonal to directions affecting the global knowledge while still enabling effective training on local datasets.
@@ 2.3 Proximal Gradient Projection
None (FedAvg) 56.13±0.78
L2 Distance 59.80±1.12 56.35±2.85 (- 3.45)
KL-Divergence 60.31±2.07 50.88±3.55 (- 9.43)
@@ 3.1 Preliminary: Overview of SAM
In the above equation, the inner maximization identifies a parameter perturbation $\epsilon$ that maximizes loss change within the $\rho$-ball neighborhood. This is practically approximated by a single re-scaled gradient step ${\boldsymbol{\epsilon}}^{*}$$=$$\rho{\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})/\|{\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})\|_{2}$. The outer minimization is then conducted by a base optimizer, such as SGD [50], taking the gradient ${\nabla}_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w}+{\boldsymbol{\epsilon}}^{*})$ at the perturbed weights. SAM demonstrates an exceptional ability to perform well across different model structures [11, 77] and tasks [66, 1] with high generalization performance. In FL, applying SAM improves the generalization of each client’s local model [55, 6]. However, since this approach only addresses local objectives, its effectiveness in generalization is mostly confined to local data distributions [60, 61] and still encounter conflicts with the proximal objective.
@@ 3.2 Adversarial Proximal Perturbation
In the above procedures, the model is updated using the gradient of the local loss $\mathcal{L}^{k}_{\mathrm{local}}$, while the proximal loss $\mathcal{L}_{{p}}^{k}$ plays only an implicit role in perturbing weights. We emphasize that unlike SAM, which employs the same loss for both weight perturbations and parameter updates, FedSOL distinguishes between the roles of these two types of losses to address the knowledge trade-off in FL. A more detailed comparison of these two distinct approaches in the FL context is discussed in Appendix G.
@@ 3.5 Theoretical Analysis
Meanwhile, Proposition 2 compares the change of local loss $\mathcal{L}^{k}_{\mathrm{local}}$ under FedSOL to its counterpart in FedAvg. This proposition suggests that, although FedSOL calculates the local gradient at perturbed weights, its impact on the local loss is almost identical to that of FedAvg. This implies that FedSOL effectively reduces local loss without significantly slowing down the learning process. As a result, FedSOL successfully mitigates the conflict between the proximal objective and the local objective. The detailed proofs of the propositions are provided in Appendix K.
@@ 4.2 Proximal Orthogonality of FedSOL
FedNova [64] 88.24${}_{(1.37)}\,$$\downarrow$ 10.00${}_{(\scriptsize{\textit{Failed}})}\,$$\downarrow$ 10.00${}_{(\scriptsize{\textit{Failed}})}\,$$\downarrow$ 64.67${}_{(0.77)}\,$$\downarrow$ 70.04${}_{(0.45)}\,$$\downarrow$ 53.07${}_{(3.30)}\,$$\downarrow$ 21.89${}_{(1.71)}\,$$\downarrow$ 38.94${}_{(2.34)}\,$$\downarrow$ 15.03${}_{(3.74)}\,$$\downarrow$
Scaffold [28] 94.18${}_{(0.32)}\,$$\downarrow$ 10.00${}_{(\scriptsize{\textit{Failed}})}\,$$\downarrow$ 10.00${}_{(\scriptsize{\textit{Failed}})}\,$$\downarrow$ 71.92${}_{(0.17)}\,$$\uparrow$ 75.49${}_{(0.21)}\,$$\uparrow$ 21.46${}_{(1.75)}\,$$\downarrow$ 16.89${}_{(2.25)}\,$$\downarrow$ 18.07${}_{(0.04)}\,$$\downarrow$ 32.04${}_{(0.07)}\,$$\downarrow$
@@ 4.3 Performance on Data Heterogeneity
Table 2 presents a comparison between our approach, FedSOL, and other baselines. The results show that many recently proposed FL methods tend to underperform even when compared to the standard FedAvg baseline. A similar observation is also reported in [75, 33], which points out that many FL methods are sensitive to the learning scenarios. In contrast, FedSOL achieves state-of-the-art results in most cases, consistently outperforming FedAvg across all evaluated scenarios. Particularly, FedSOL shows remarkable improvement at high heterogeneity levels, such as in Sharding ($s$=2) and LDA ($\alpha$=0.05) scenarios. We further provide the learning curves in Appendix C, comparison to other weights perturbation strategies in Appendix H, personalized FL performance in Appendix D, and performance on larger datasets in Appendix E.
@@
We examine the learning factors that influence FedSOL’s performance in Figure 4. Note that the sampling of clients simulates communication failures with inactive devices. In our experiments, FedSOL consistently surpasses FedAvg across various factors, achieving its best performance within the $\rho$ range between 0.5 and 2.0. Most notably, FedSOL’s gains increase as a smaller portion of clients participate in each round. For example, FedAvg’s performance significantly declines at a sampling ratio of 0.02, falling to near-random accuracy. However, FedSOL remains robust under such conditions. Further comparisons are provided in Appendix J.
@@
Vgg11 FedAvg 41.30±1.07 50.02±4.25 61.79±9.88
FedProx 40.45±1.41 31.07±6.72 63.47±2.68
FedNTD 60.55±2.14 56.62±2.64 69.82±2.27
FedSOL 56.39±1.40 74.74±0.04 78.38±1.12
Res18 FedAvg 49.92±0.62 76.98±2.90 57.91±1.27
FedProx 59.00±2.58 82.09±2.35 75.84±1.58
FedNTD 57.79±3.42 78.50±0.18 76.87±0.57
FedSOL 66.32±0.48 85.97±0.04 80.59±0.11
SL-ViT FedAvg 35.48±2.09 53.94±5.17 72.44±1.91
FedProx 38.73±1.23 58.25±4.23 74.10±1.23
FedNTD 47.59±2.84 61.46±1.76 71.65±1.71
FedSOL 47.95±1.51 67.19±0.33 77.96±0.47
@@ 4.4 Ablation Study
The results in Table 5 reveal that perturbing only the last classifier layer (Head in Table 5) is sufficient for FedSOL. The performance reaches as high as the full-model perturbation, yet the required computation is considerably lower. Interestingly, perturbing all layers except the head (Body in Table 5) incurs nearly the same computational cost yet results in diminished performance, highlighting the importance of the later layers. We conduct further experiments on larger models in Appendix F and discuss the local computational cost in Appendix I.
@@ 6 Related Work
In CL, many approaches have adopted orthogonal learning, which align new task gradients orthogonal to old task loss spaces [13, 56, 19, 41]. This typically involves retaining data or gradients as memory [9, 10]. In FL, few recent attempts have applied similar strategies [4, 42]. For example, FOT [4] uses a random Gaussian matrix with SVD, and GradMA [42] employs gradient memory and quadratic programming, both requiring substantial computational resources. In our work, we use the proximal objective to preserve knowledge and update using a local gradient orthogonal to it. However, we observe that directly opposing the proximal gradient can adversely affect performance, a finding that aligns with recent CL research about the drawbacks of direct gradient projection [76, 70]. To address this, we propose FedSOL, which implicitly finds the orthogonal local gradient.
@@ Appendix A Table of Notations
$k$ Index for clients ($k\in[K]$)
$g$ Index for global server
@@
$\mathcal{D}$ Whole dataset
$\mathcal{D}^{k}$ Local dataset of the $k$-th client
$\alpha$ Concentration parameter for the Dirichlet distribution
$s$ The number of shards per user
@@
$\beta,\mu$ Multiplicative coefficient for the proximal loss
$\gamma$ Learning rate
$\tau$ Temperature to be divided in the softmax probability distribution
$\rho$ Perturbation Strength for SAM-related algorithms
$\boldsymbol{\Lambda}$ Vector consists with scaling parameters for perturbation vector in SAM-related algorithms
@@
$\boldsymbol{w}_{g}$ Weight of the global server model
$\boldsymbol{w}_{k}$ Weight of the $k$-th client model
$\|\boldsymbol{w}_{g}-\boldsymbol{w}_{k}\|$ Collection of $L^{2}$-norm between server and client models, among all rounds.
@@
$\mathcal{L}^{k}_{\mathrm{local}}$ Local objective for the $k$-th client
$\mathcal{L}^{k}_{p}$ Proximal objective for the $k$-th client
@@ B.2 Datasets
  • CINIC-10 [12] (90,000 / 90,000): is a combination of CIFAR and downsized ImageNet [14], which is compiled to serve as a bridge between the two datasets. The data is augmented using Random Cropping, Random Horizontal Flipping, and Normalization.
  • PathMNIST [71] (110,000 / 7,180): contains non-overlapping patches from Hematoxylin & Eosin stained colorectal histology slide images. The data is augmented using Random Horizontal Flipping, and Normalization.
  • TissueMNIST [71] (189,106 / 47,280): contains microscope images of human kidney cortex cells, which are segmented from 3 reference tissue specimens. The data is augmented using Random Horizontal Flipping, and Normalization. The data is converted to RGB images.
@@ Appendix D Personalized Performance
Appendix E Performance on Larger Datasets
In Table 2, we show that FedSOL consistently achieves performance gains, while most existing FL (Federated Learning) methods are sensitive to learning setups. Although FedNTD marginally outperforms FedSOL in some CIFAR-10 cases, it significantly falls behind in most others. In Table 10, we conducted experiments on the CIFAR-100 and ImageNet-100 datasets. We used the ResNet-18 model, distributing each dataset across 100 clients with a sampling ratio of 0.1 and optimized for 5 local epochs. Our observations indicate that FedSOL maintains its effectiveness in both datasets, whereas FedNTD’s performance decreases in CIFAR-100.
Table 10: Test Accurac on CIFAR-100 and ImageNet-100.
Method CIFAR-100 ImageNet-100
s = 5 s = 10 $\alpha$ = 0.05 $\alpha$ = 0.1
FedAvg 42.43 53.23 48.69 43.41
FedProx 39.03 $\downarrow$ 48.38 $\downarrow$ 46.75 $\downarrow$ 34.49 $\downarrow$
FedNTD 39.32 $\downarrow$ 52.23 $\downarrow$ 48.35 $\downarrow$ 44.08 $\uparrow$
FedSOL 44.21$\uparrow$ 53.78$\uparrow$ 49.25$\uparrow$ 44.97$\uparrow$
Appendix F Head Perturbation in Larger Models
To validate the effectiveness of partial perturbation strategy, we extended the comparison experiment in Table 5 to Table 11. The results on VggNet-11 and ResNet-18 indicate that perturbing only the last classifier layer (head) is almost as effective as perturbing the entire model (full), saving significant computational cost.
Table 11: Effect of partial perturbation on CIFAR-10 ($\alpha$=0.1).
Model FedAvg FedSOL (full) FedSOL (head)
VggNet-11 41.30 56.44 (+ 15.14) 56.39 (+ 15.09)
ResNet-18 49.92 66.69 (+ 16.77) 66.32 (+ 16.04)
Appendix G Comparison to the Sharpness-Aware Optimization
G.1 SAM Optimization in FL
@@ G.2 Limitations of using SAM in FL
Figure 12: Performance on CIFAR-10 ($\alpha$=0.1) datasets.
Appendix H Other Perturbation Strategies
@@
We exclude the strategies that obtaining proximal loss at the perturbed weights using the local loss gradient i.e., $\mathcal{L}_{p}(\boldsymbol{w}_{k}+\boldsymbol{\epsilon}_{c}^{*}$), where $\boldsymbol{\epsilon}_{c}^{*}$ $=$ $\rho\,\frac{\boldsymbol{g}_{p}+\boldsymbol{g}_{l}}{\lVert\boldsymbol{g}_{p}+\boldsymbol{g}_{l}\rVert}$, as it leads the learning to diverge. The detailed formulation for each method is provided in Table 12 with its corresponding performance. The results in Table 12 demonstrates that utilizing the local loss gradient at weights perturbed by the proximal loss gradient ($A_{1}$ in Table 12) yields outperforms the other approaches. We suggest that our FedSOL is an effective way to integrate proximal restriction effect into SAM optimization in FL.
@@
Appendix I Client-side Computational Cost
In Table 5, we analyze the FLOPS required for FedSOL and note that perturbing only the head part is almost as effective as full perturbation, yet it requires only 33% more computation than FedAvg. Although FedSOL necessitates backward computation twice, this does not lead to increased GPU memory usage, as FedSOL does not store these gradients simultaneously. The slight increase in memory usage ($<0.01$%) arises from calculating the adaptive perturbation strength for each layer. In Table 13, we present the latency for a single local step with the ImageNet-100 dataset, using a ResNet-18 model on an NVIDIA A6000 GPU. We measured FedSOL’s latency using an L2 proximal loss for a fair comparison. While FedSOL involves additional local computation, we emphasize that in FL, the energy consumed for communication typically surpasses that for computation. The faster convergence of FedSOL substantially reduces the total energy consumption. In the ImageNet-100 experiment detailed in Table 10, for example, FedSOL achieves FedAvg’s 300-round performance by round 238.
Table 13: Measured latency for a single local step.
Usage FedAvg FedProx FedNTD FedSAM FedSOL
Latency 2.846 s 2.865 s 2.996 s 3.380 s 2.900 s
Appendix J Effect of Learning Factors on FedSAM
Figure 12 presents the impact of learning factors on FedSAM [55] and FedASAM [6], showing that both methods are more sensitive to these factors compared to FedSOL. In most of our experiments, SAM-related FL methods shows inferior performance compared to FedAvg. This may be because SAM, aiming to enhance local model generality, becomes less effective for small models, under conditions of high data heterogeneity, or with a low sampling ratio. Future research is expected to identify the conditions where employing SAM on the local side becomes beneficial.
Appendix K Proof of Proposition